in an ap the first term is 1, if second term, tenth term and thirty-fourth terms forms a GP, then what is the fourth term of the GP series? 

Grade:12

2 Answers

Vikas TU
7 years ago
t1 = 1
t2 = 1+d
t10 = 1 + 9d
t34 = 1 + 33d
these are all in G.P. then,
t2/t1 = t34/t10
solve for d and thewrite t4 = 1 + 3d thus.
krunal
7 years ago
here, a1=1 and 2,10,34 are  forming gp. 
so we can write a2= a+d, a10= a+9d, and a34= a+33d.
 now its given that they are forming gp so by property we can write,
(a+9d)square= (a+d)(a+33d).      (a=1)
on solving we get d=1/3.
so 4th term will be = a+3d= 2.